### v2

@@ 5.1.3 Parameter settings
For all methods, we tune the regularization parameters via grid search over $\{10^{-2},10^{-1},1,10,10^{2}\}$. For graph-based methods, we set the number of k-nearest neighbors to 5. For methods utilizing $\ell_{2,p}$-norm, we set $p=1$. For STPCA-MP, we set $\mathbf{M}=\mathbf{I}$ in most experiments except the one that discusses the choice of $\mathbf{M}$. The direction set $D^{o}$ is also tuned by a grid-search strategy in the range $\{\{1,3,2\},\{2,3,1\}\}$. The other parameters remain the default in the open code provided by the authors. The best results of all methods are presented.
@@ 5.1.4 Evaluation methodology and metrics
In clustering experiments, after a UFS method has finished selecting features, we evaluate its performance by first conducting K-means clustering and then mapping the obtained pseudo labels to the real labels by adopting the Kuhn-Munkres algorithm [41]. Due to the K-means’ dependence on initialization, we repeat the clustering 30 times. The clustering metrics we use are as follows:
@@
    Accuracy (ACC): ACC refers to the clustering accuracy and is computed according to (35) and (36), where $n$ is the total number of samples, $map(i)$ denotes the clustering label of the $i$-th sample after mapping to the real label, and $label(i)$ represents the ground truth label. A larger ACC indicates better performance.
    $\text{ACC}=\frac{1}{n}\sum\limits_{i=1}^{n}\delta\left(map(i),label(i)\right).$ (35)
    $\delta(p,q)=\left\{\begin{aligned} &1\quad\text{if}\;p=q,\\ &0\quad\text{otherwise}.\end{aligned}\right.$ (36)
@@
    Normalized Mutual Information (NMI) [42]: NMI is utilized to describe the mutual dependence between clustering results (after being mapped into the real labels) and ground truth labels. Given the clustering results $\mathbf{m}$ and the ground truth labels $\mathbf{l}$ (both including all samples), $I(\mathbf{m},\mathbf{l})$ denotes the mutual information between $\mathbf{m}$ and $\mathbf{l}$. $H(\mathbf{m})$ and $H(\mathbf{l})$ are the entropy of $\mathbf{m}$ and $\mathbf{l}$ respectively. The same as ACC, a larger NMI means better clustering performance.
@@
For orientation-specific data such as multi-channel time series, the temporal dependency of the elements within the same channel makes the clustering metrics unreliable due to the feature independence hypothesis in K-means. Therefore, we define a metric as follows:
@@
    $\text{POC}=c/h,$ (38)
    where $c$ is the number of correctly selected features, and $h$ is the total number of selected features at each execution. POC reflects a UFS method’s ability to select discriminative features. In the experiments, we select features with maximal between-class variance (BCV) as the correct ones. We compute POC on slice-wise data instead of ACC and NMI.
@@ 5.2 Datasets
Before detailing specific datasets, we first categorize tensor datasets based on their structural organization. As shown in Figure 2, third-order tensor datasets can be divided into two scenarios: tube-wise and slice-wise, according to their organization. Slice-wise data is orientation-specific and exhibits feature relations along the 1-mode, while tube-wise data exhibits feature relations along both the 1-mode and 2-mode.. This distinction provides a basis for evaluating the ability of UFS methods to utilize tensor structure information. In our evaluation, we compare clustering ACC and NMI on tube-wise data and POC on slice-wise data. Experimental results demonstrate the effectiveness of the proposed method in both scenarios.
Table 2: Two scenarios for third-order data tensors. Assume each frontal slice is a sample. Different colors refer to different variables. This paper regards variables as features. Tube-wise scenario: elements within one fiber belong to one variable. Slice-wise scenario: elements within one horizontal slice belong to one variable.
@@
    Orbit: As an example of slice-wise data, Orbit comprises 3-dimensional orbit samples with two distinct radii. Each sample includes three channels representing orbit coordinates, with the remaining channels containing random Gaussian noise.
@@
    Array: As an example of tube-wise data, Array is constructed using a $10\times 10$ synthetic radar unit array that receives a series of echoes from a stationary target. Each echo is represented as a matrix of unit sampling values, treated as a sample. Random measurement errors are introduced to half of the echo matrices to create two distinct classes.
We evaluate our method on nine real-world datasets, including: 1) six tube-wise datasets: COIL20 [54], USPS [52], BreastMNIST [44], Imm40 [51], PIE [13] and JAFFE [43]; 2) four slice-wise datasets: UCIHAR [45] and UCIDSA [46], UMCS [^1] https://www.kaggle.com/c/seizure-detection/data and ASCH [53]. The detailed information of each dataset is shown in Table 3. PIE, Imm40, and JAFFE contain frontal facial images, labeled according to individuals. BreastMNIST is a medical dataset containing breast ultrasound images, from which we construct the data tensor using the test set. USPS is a dataset of handwritten digits. UCIHAR and UCIDSA are human activity datasets collected from multiple subjects using motion sensors. UCIDSA contains multi-channel signals, while UCIHAR comprises extracted statistical features (e.g. mean, std, max, entropy). UMCS and ASCH are electroencephalogram (EEG) datasets collected for disease research, capturing brain activity signals. Due to the large volume of chosen slice-wise data, we retrieve a subset of samples to construct datasets. For UCIDSA, we extract the first $125$ sample points from all $45$ channels covering $45$ activities. For UCIHAR, we select $9$ individuals with all $561$ features sampled during $6$ activities. For UMCS, we collect the first $125$ sample points in all $65$ channels of the first human patient, classified into ictal and interictal segments. For ASCH, we collect the first $640$ sample points in all $16$ channels from healthy groups and groups with symptoms. All datasets are normalized and bounded within $[-1,1]$ for consistency.
@@ 5.3 Synthetic experiment
In this experiment, we conduct four typical UFS methods on synthetic datasets with the grid search strategy. The dataset formulation and the best experimental results of comparative methods are illustrated in Figure 3. On Orbit, we retrieve three top-ranked features in each method’s result and plot their samples in a subset of data. If a method selects the correct features (i.e. the three orbit coordinates), then the scatter plot will contain rings of two different radii. Otherwise, it will be difficult to distinguish the two classes. On Array, we retrieve the feature score map corresponding to the best POC for each method, analyzing the match degree between the top-ranked feature positions and the units with measurement error. We conclude the observations as follows.
@@
    Tensor-based methods are generally more capable of analyzing the data structure than non-tensor-based methods. Both CPUFS and STPCA-MP select the orbit coordinates on Orbit. Although CPUFS fails to precisely locate the units with error on Array, its score map clearly shows a grid-like characteristic that connects the rows and columns to analyze feature relations, capturing a roughly correct attention region. A similar pattern also appears in STPCA-MP, where the column-wise relations are learned. In contrast, RNE and NLGMS isolate the features, losing the tensor structure information. Especially in RNE, the scores on Array seem to be randomly distributed.
@@
    Our proposed method performs the best in selecting discriminative features on the two synthetic datasets. STPCA-MP selects the correct features on both datasets. In particular, the score map on Array is relatively sparse in rows and columns. By taking advantage of data orientation, STPCA-MP can more easily mine the latent feature relations and locate discriminative features.
@@ 5.4 Clustering experiment on tube-wise data
In this experiment, we perform clustering on five tube-wise real-world datasets. The regularization parameters of all methods go through a grid search in $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$. We set the selected features to $\{50,100,150,200,250,300\}$ on COIL20, JAFFE, BreastMNIST , Imm40, and $\{10,30,50,70,90,110\}$ on USPS. The best results and the corresponding training time of all methods are given in Table 4. The detailed ACC and NMI curves are illustrated in Figure 4. By analyzing the results, we come out with the following conclusions:
@@
    Features selection improves data compactness. On each dataset, one or more methods are capable of achieving performance that approaches or even surpasses the baseline while utilizing only a subset of features. In particular, the proposed STPCA-MP achieves significantly higher clustering metrics than all features on Jaffe, BreastMNIST, and Imm40.
@@
    STPCA-MP delivers reliable and outstanding performance in clustering tasks. Firstly, compared to graph-based methods like SOGFS, STPCA-MP obtains more global data information by exploiting the reconstruction term, which guarantees the basic clustering performance. We can see the curves of STPCA-MP, on the whole, have a smaller oscillation in amplitude and share better consistency in trend between ACC and NMI. Besides, STPCA-MP reflects a more natural decreasing (or increasing) progression in variation pattern. Secondly, the T-SVDM-based sparse framework is beneficial for extracting tensor structure information, enabling STPCA-MP to describe the feature correlation in a more delicate manner, compared to CPD-based tensor methods (CPUFS and CPUFSnn), Tucker-based tensor method (MSPCA), and non-tensor-based methods (SPCAFS, RNE). Therefore, STPCA-MP demonstrates superior performance, consistently ranking at the forefront on all datasets.
@@
    STPCA-MP achieves the best performance at a smaller training time cost. On all datasets in Table 4, STPCA-MP takes less than $1$ second to complete training and feature selection, due to its slice-by-slice operation. Note that this is accomplished with reasonable prior knowledge of data organization. As a comparison, NLGMS, also showing remarkable average performance, brings much more computational cost as it overlooks the tensor structure by vectorization operation.
Table 4: Metrics of the comparative methods on real-world datasets. The best ACC, best NMI, and mean POC regarding grid-search range are shown. The training time corresponding to the best result is recorded in seconds. The best result on each dataset is boldfaced, while the second best one is underlined and italic.
@@
Figure 4: The ACC and NMI curves on tube-wise datasets. The upper row shows the ACC curves, while the bottom row displays the NMI curves.
@@ 5.5 POC comparison experiment on slice-wise data
For slice-wise datasets, we view each channel (row vector) as a feature. Therefore, we perform this experiment in the following steps: First, we implement a UFS algorithm and obtain the scores for each element. Second, we sum row-wise scores to evaluate features, using POC to measure the agreement between top-ranked features and those with maximal between-class variance (BCV). We compute BCV after the data are zero-centered.
The BCV of selected features on four slice-wise datasets is shown in Figure 5, where the results corresponding to the greatest POC of typical comparative methods are given. The mean POC of all methods is illustrated in Table 4. We have the following observations:
@@
    For each dataset, there exists a small set of features with prominent BCV, judging by the peaks on the BCV curves in Figure 5. Greater BCV usually implies better inter-class discriminability.
@@
    SPCA-based methods are generally more capable of selecting the features of the greatest BCVs. SPCAFS, MSPCA, and STPCA-MP have a stabler performance in terms of POC, owing to the encoder-decoder framework that minimizes the reconstruction error as well as maximizes the variance among samples on the global manifold. Attributed to the sparse tensor decomposition, MSPCA also achieves relatively stable performance on slice-wise data, referring to its mean POC in Table 4.
@@
    STPCA-MP has the advantage of handling slice-wise data over comparative methods. Due to the exploitation of data orientation, STPCA-MP can better learn the intrinsic relationship between different channels in a slice-wise tensor, selecting a considerable amount of BCV peaks in Figure 5. Although the best result of NLGMS is compatible to our method on some datasets, STPCA-MP demonstrates enhanced stability in selecting discriminative features, according to the mean POC regarding grid-search range in Table 4. On average, STPCA-MP is more effective toward slice-wise data, compared to the state-of-the-arts.

### v3

@@ 5.1.3 Parameter settings
All methods employ consistent experimental protocols: (1) Regularization parameters are determined through grid search over $\{10^{-2},10^{-1},1,10,10^{2}\}$; (2) Uniform feature dimensionality is maintained across comparative methods. Specific configurations include: (i) Graph-based methods use $k=5$ nearest neighbors; (ii) $\ell_{2,p}$-norm methods fix $p=1$; (iii) STPCA-MP initializes $\mathbf{M}=\mathbf{I}$ for standard experiments (excluding the dedicated $\mathbf{M}$ analysis); (iv) Direction set $D^{o}$ undergoes grid search in $\{{1,3,2\},\{2,3,1\}}$. Remaining parameters follow default settings from authors’ original implementations. All reported results reflect optimal parameter configurations.
@@ 5.1.4 Evaluation methodology and metrics
Following feature selection by each UFS method, we assess clustering performance through a standardized evaluation pipeline:
  1. [1)]
  2. 1.
    Cluster Generation: Apply $K$-means clustering to the selected features.
  3. 2.
    Label Alignment: Establish correspondence between pseudo-labels and ground truth using the Kuhn-Munkres algorithm [44].
  4. 3.
    Stability Control: Repeat the clustering process 30 independent times and average the result to account for $K$-means’ initialization sensitivity.
The following clustering metrics are employed for quantitative evaluation:
@@
    Accuracy (ACC): The clustering accuracy (ACC) measures the agreement between cluster assignments and ground truth labels, computed as:
    $\mathrm{ACC}=\frac{1}{n}\sum_{i=1}^{n}\delta(\mathrm{map}(i),\mathrm{label}(i))$ (39)
    where $\delta(p,q)$ is the indicator function defined in (40), $n$ denotes the total sample size, $\mathrm{map}(i)$ represents the mapped cluster label of the $i$-th sample, and $\mathrm{label}(i)$ indicates the ground truth label. Higher ACC values correspond to better clustering performance. The indicator function is given by:
    $\delta(p,q)=\begin{cases}1&\text{if }p=q\\ 0&\text{otherwise}\end{cases}$ (40)
@@
    Normalized Mutual Information (NMI) [45]: NMI is utilized to describe the mutual dependence between mapped clustering labels and ground truth labels. Given the clustering results $\mathbf{m}$ and the ground truth labels $\mathbf{l}$ (both including all samples), $I(\mathbf{m},\mathbf{l})$ denotes the mutual information between $\mathbf{m}$ and $\mathbf{l}$. $H(\mathbf{m})$ and $H(\mathbf{l})$ are the entropy of $\mathbf{m}$ and $\mathbf{l}$ respectively. A larger NMI means better clustering performance.
@@
    Silhouette score (SS) [16]: SS is a widely-used metric for evaluating clustering quality that measures both intra-cluster compactness and inter-cluster separation. For a given sample $i$, the silhouette score $s(i)$ is computed as:
    $s(i)=\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$ (42)
    where $a(i)$ is the average distance between sample $i$ and all other points in the same cluster (intra-cluster distance), and $b(i)$ is the smallest average distance between $i$ and points in any other cluster (neighboring cluster distance). The global silhouette score SS for the entire dataset ranges from $-1$ to $1$, with values closer to $1$ indicating better-defined clusters:
    $\text{SS}=\frac{1}{N}\sum_{i=1}^{N}s(i).$ (43)
  5. 4.
@@
For orientation-specific data (e.g., multi-channel time series), the temporal dependencies within channels violate the feature independence assumption underlying $K$-means clustering. This fundamental mismatch renders standard clustering metrics unreliable. We therefore propose the Proportion of Correctly selected features (POC) as an alternative evaluation metric:
@@
    $\text{POC}=\frac{c}{h},$ (44)
    where $c$ denotes the number of features selected by both the unsupervised feature selection (UFS) method and the maximal between-class variance (BCV) criterion, and $h$ represents the total features selected per execution. Suppose data has been centralized, the BCV for the $(i,j)$-th element in a frontal slice is defined as:
    $\text{BCV}(i,j)=\frac{1}{K}\sum_{k=1}^{K}\left(\mu^{(k)}_{ij}-\mu_{ij}\right)^{2},\quad i=1,\dots,d_{1};\ j=1,\dots,d_{2}$ (45)
    where $\mu_{ij}$ is the sample mean of the $(i,j)$-th element across all samples, $\mu^{(k)}_{ij}$ is the sample mean within class $k$, and $K$ is the number of classes. Here, $d_{1}$ and $d_{2}$ denote the dimensionalities of the first and second modes, respectively. For slice-wise data where each channel is treated as a feature (See Table 5 and Section 5.2), the BCV for the $f$-th channel is obtained by aggregating along the temporal dimension:
    $\text{BCV}_{f}=\sum_{j=1}^{d_{2}}\text{BCV}(i,j)\quad\text{for fixed channel index }i=f.$ (46)
    This formulation quantifies the cumulative class-discriminative power of the entire time series in channel $f$. The POC metric thereby directly evaluates a method’s capability to identify discriminative features where conventional clustering metrics fail.
@@ 5.2 Datasets
Before delving into specific datasets, we first categorize tensor datasets based on their structural organization. As illustrated in Table 5, third-order tensor datasets can be classified into two scenarios: tube-wise and slice-wise, according to their organizational patterns. Slice-wise data is orientation-specific, exhibiting feature correlations along the 1-mode, whereas tube-wise data demonstrates feature correlations along both the 1-mode and 2-mode. This distinction serves as a foundation for evaluating the capability of UFS methods to leverage tensor structural information. In our evaluation, we compare ACC, NMI, and SS for tube-wise data, and use the POC for slice-wise data. Experimental results validate the effectiveness of the proposed method in both scenarios.
Table 5: Two scenarios for third-order data tensors. Each frontal slice corresponds to a distinct sample. Variables (represented by distinct colors) are treated as features. This study adopts two feature interpretation scenarios. Tube-wise: Elements within a single fiber constitute one feature variable. Slice-wise: Elements within a horizontal slice comprise one feature variable.
@@
    Orbit (Figure 4(a)): As an example of slice-wise data, Orbit comprises 3-dimensional orbit samples with two distinct radii. Each sample includes three channels representing orbit coordinates, with the remaining channels containing random Gaussian noise.
@@
    Array (Figure 4(f)): As an example of tube-wise data, Array is constructed using a $10\times 10$ synthetic radar unit array that receives a series of echoes from a stationary target. Each echo is represented as a matrix of unit sampling values, treated as a sample. Random measurement errors are introduced to half of the echo matrices to create two distinct classes.
We evaluate our method on nine real-world datasets, including: 1) six tube-wise datasets: COIL20 [57], USPS [55], BreastMNIST [47], Imm40 [54], PIE [13] and JAFFE [46]; 2) four slice-wise datasets: UCIHAR [48] and UCIDSA [49], UMCS [^8] https://www.kaggle.com/c/seizure-detection/data, and ASCH [56]. The detailed information of each dataset is shown in Table 6. PIE, Imm40, and JAFFE contain frontal facial images, labeled according to individuals. COIL20 comprises images of various objects captured from multiple viewing angles. BreastMNIST is a medical dataset containing breast ultrasound images, from which we construct the data tensor using the test set. USPS is a dataset of handwritten digits. UCIHAR and UCIDSA are human activity datasets collected from multiple subjects using motion sensors. UCIDSA contains multi-channel signals, while UCIHAR comprises extracted statistical features (e.g. mean, std, max, entropy). UMCS and ASCH are electroencephalogram (EEG) datasets collected for disease research, capturing brain activity signals. Due to the large volume of chosen slice-wise data, we retrieve a subset of samples to construct datasets. For UCIDSA, we extract the first $125$ sample points from all $45$ channels covering $19$ activities. For UCIHAR, we select $9$ individuals with all $561$ features sampled during $6$ activities. For UMCS, we collect the first $125$ sample points in all $65$ channels of the first human patient, classified into ictal and interictal segments. For ASCH, we collect the first $640$ sample points in all $16$ channels from healthy groups and groups with symptoms. All datasets are normalized and bounded within $[-1,1]$ for consistency. All the above data preprocessing is consistent across all comparative methods, including ours.
@@ 5.3 Synthetic experiment
In this experiment, we apply four typical UFS methods to synthetic datasets using the grid search strategy. The dataset construction and the optimal experimental results of the comparative methods are presented in Figure 4.
For the Orbit dataset (Figure 4(a)), we retrieve the three top-ranked features from the results of each method and plot their corresponding samples within a data subset. If a method successfully selects the correct features (i.e., the three orbit coordinates), the scatter plot will display rings with two distinct radii. Conversely, when the wrong features are chosen, distinguishing between the two classes becomes challenging.
Regarding the Array dataset (Figure 4(f)), we retrieve the feature score map associated with the best POC for each method. Then, we analyze the degree of match between the positions of the top-ranked features and the units with measurement errors.
Considering the experimental results on both synthetic datasets, the key observations are summarized as follows:
@@
    Tensor-based methods generally outperform non-tensor-based methods in analyzing data structure. Both CPUFS and STPCA-MP correctly identify the orbit coordinates on the Orbit dataset (Figure 4(d)(e)). Although CPUFS fails to precisely pinpoint the error units on the Array dataset, its score map exhibits a distinct grid-like pattern (Figure 4(i)), which captures the row-column feature relationships and identifies an approximately correct region of interest. Similarly, STPCA-MP learns the column-wise relationships (Figure 4(j)). In contrast, RNE and NLGMS treat features in isolation, thus losing the valuable tensor structure information. Notably, for RNE in Figure 4(g), the scores are seemingly randomly distributed.
@@
    Our proposed method excels in selecting discriminative features on both synthetic datasets. STPCA-MP selects the correct features on both datasets (Figure 4(e)(j)). Specifically, the score map for the Array dataset (Figure 4(j)) shows relatively sparse rows and columns. By leveraging the data’s orientation, STPCA-MP can more effectively uncover latent feature relationships and identify discriminative features.
@@ 5.4 Clustering experiment on tube-wise data
In this experiment, we perform clustering on five tube-wise real-world datasets. The regularization parameters of all methods go through a grid search in $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$. We set the selected features to $\{50,100,150,200,250,300\}$ on COIL20, JAFFE, BreastMNIST , Imm40, and $\{10,30,50,70,90,110\}$ on USPS. The best results and the corresponding training time of all methods are given in Table 7. The detailed ACC, NMI, and SS curves are illustrated in Figure 5. From these results, we draw three main conclusions:
@@
    Features selection enhances data compactness. Multiple methods achieve performance comparable to or superior to the baseline while utilizing only a subset of features. Notably, STPCA-MP significantly outperforms All Features on PIE (Figure 5(a)(g)(m)), Jaffe (Figure 5(c)(i)(o)), BreastMNIST (Figure 5(d)(j)(p)), and Imm40 (Figure 5(e)(k)(q)).
@@
    STPCA-MP demonstrates reliable and superior clustering performance. The proposed method captures more global data information through its reconstruction term, outperforming graph-based approaches like SOGFS, CPUFS, and CPUFSnn. This is evidenced by: (i) smoother performance curves with smaller amplitude oscillations (e.g. Figure 5(e)(k)(q)), and (ii) more consistent trends across ACC, NMI, and SS metrics (Figure 5(c)(i)(o)). Furthermore, STPCA-MP’s T-SVDM-based framework effectively extracts tensor structural information, enabling more precise feature correlation modeling than: (a) CPD-based methods (CPUFS, CPUFSnn), (b) Tucker-based MSPCA, and (c) non-tensor methods (SPCAFS, RNE) on datasets with relatively strict feature spatial distributions (PIE, JAFFE, and Imm40). On datasets with irregular feature space distributions (COIL20 and USPS), STPCA-MP does not demonstrate a significant advantage in average performance, yet maintains top-ranked results across all tested feature selection dimensions (Figure 5(b)(h)(n)) and ultimately attains the optimal results (Table 7).
@@
    STPCA-MP combines high performance with computational efficiency. As shown in Table 7, STPCA-MP completes training and feature selection in under 1 second per dataset through its slice-by-slice operations. This efficiency is achieved while maintaining proper data structure awareness. In contrast, NLGMS and SPCAFS incur substantially higher computational costs due to their vectorization approach, despite showing competitive accuracy.
[…]
@@
Figure 5: The ACC, NMI, and SS curves on tube-wise datasets. The upper, middle, and bottom row displays the ACC, NMI, and SS curves, respectively.
@@ 5.5 POC comparison experiment on slice-wise data
For slice-wise datasets where each channel (row vector) is treated as a feature, the experiment proceeds through two key phases: (i) Feature Scoring: executing the UFS algorithm to generate element-wise importance scores, followed by (ii) Feature Evaluation: aggregating scores row-wise for channel-level rankings, computing BCV on zero-centered data, and computing POC by comparing top-ranked features with maximal BCV features. The number of selected features is set to $15$, $100$, $15$, and $6$ on UCIDSA, UCIHAR, UMCS, and ASCH, respectively.
The BCV distribution of selected features across four slice-wise datasets is visualized in Figure 6, displaying results corresponding to each method’s maximal POC performance. Mean POC values are quantitatively compared in Table 7. We have the following observations:
@@
    Each dataset exhibits a subset of features with significantly higher BCV, as evidenced by distinct peaks in Figure 6(a)(g)(m)(s). These peaks correlate strongly with inter-class discriminative power.
@@
    SPCA-based methods demonstrate superior BCV feature selection capability. SPCAFS, MSPCA, and STPCA-MP maintain stable POC performance through their shared encoder-decoder framework, which simultaneously minimizes reconstruction error and maximizes global sample variance. Notably, MSPCA’s sparse tensor decomposition contributes to its competitive performance on slice-wise data, as confirmed by its mean POC in Table 7.
@@
    STPCA-MP outperforms competitors in slice-wise data processing. By explicitly modeling data orientation, STPCA-MP effectively captures cross-channel relationships within slice-wise tensors, consistently selecting BCV peaks, as shown in Figure 6(b)(h)(n)(t). While NLGMS achieves comparable peak performance on certain datasets, STPCA-MP shows superior stability in discriminative feature selection across the full parameter range, as illustrated in Table 7, establishing it as the most effective method for slice-wise data analysis.